Use Hückel theory to write the secular determinant and determine the energies of the pi orbitals and total energy of the pi system for hexatriene
Answer #1

The Secular determinant of hexatriene is . ; L-E ER. * CÔ o | 2- 0 0 0 0 L-BE B o oo ß L-PF ß oo To o ß X-e ßo To o o ß d-BEI 11:# Energy - - - - - - - - - - - - - - I 76 V, Fi = 2 1.86287 . On solving Eg = fat 1.24787 Ez = 12 +0.44587 E4 = 2-0.445B
Answer #2
what about the equation of the determinant for the benzene in the above matrix
source: CH 312 quantum chemistry
answered by: kincade Ingram